Let "B = \begin{pmatrix}
 1 & 0 \\
 2 & 3
\end{pmatrix}"

What is B-1?

1
Expert's answer
2021-07-01T16:10:38-0400

Given "\begin{pmatrix}
 1& 0 \\
 2 & 3
\end{pmatrix}"

"A=\begin{pmatrix}
 a & b \\
 c & d
\end{pmatrix}"

"A^{-1}=\frac{1}{ad-bc}\begin{pmatrix}
 d &- b \\
 - c & a
\end{pmatrix}"

"B^{-1}=\frac{1}{1×3-2×0}\begin{pmatrix}
 3 & 0 \\
 -2 & 1
\end{pmatrix}"


"B^{-1}=\frac{1}{3}\begin{pmatrix}
 3 & 0\\
 -2 & 1
\end{pmatrix}"


"B^{-1}=\begin{pmatrix}
 1 & 0 \\
 \frac{-2}{3} & \frac{1}{3}
\end{pmatrix}"